To solve for b in this formula, the first step is to?
A= A+B+C/3
: To solve for b in this formula, the first step is to?
A= A+B+C/3

Answer by nerdybill(1052)

Solving for B means to isolate it to one side of the equation, doesn't matter which side.
.
A= A+B+C/3
Start by subtracting A from both sides:
0 = B+C/3
.
Next, subtract C/3 from both sides:
-C/3 = B